Analysis of coupled systems of kinetic equations and conservation laws

Moulay Tidriri Iowa State University

We give a survey of recent progress on the rigorous derivation and analysis of the coupling of kinetic equations and their hydrodynamic conservation laws limits for some BGK-type models.